OpenStudy (anonymous):

Write the equation of the line parallel to the like 4x-6y=-1 that intersects the function 3x-2y=12 at its x-intercept.

OpenStudy (anonymous):

slope=2/3 x=4 y=0 yi-0=2/3(0-4) yi=-8/3 y=2/3x-8/3 3y-2x=-8

OpenStudy (anonymous):

i think thats right

OpenStudy (anonymous):

hey are u there?

OpenStudy (anonymous):

yes and it is better than anything i came up with, thanks

OpenStudy (anonymous):

np